# OmniHuman-1: Rethinking the Scaling-Up of One-Stage Conditioned Human Animation Models

### v2

@@ Abstract
End-to-end human animation, such as audio-driven talking human generation, has undergone notable advancements in the recent few years. However, existing methods still struggle to scale up as large general video generation models, limiting their potential in real applications. In this paper, we propose OmniHuman, a Diffusion Transformer-based framework that scales up data by mixing motion-related conditions into the training phase. To this end, we introduce two training principles for these mixed conditions, along with the corresponding model architecture and inference strategy. These designs enable OmniHuman to fully leverage data-driven motion generation, ultimately achieving highly realistic human video generation. More importantly, OmniHuman supports various portrait contents (face close-up, portrait, half-body, full-body), supports both talking and singing, handles human-object interactions and challenging body poses, and accommodates different image styles. Compared to existing end-to-end audio-driven methods, OmniHuman not only produces more realistic videos, but also offers greater flexibility in inputs. It also supports multiple driving modalities (audio-driven, video-driven and combined driving signals). Video samples are provided on the project page.
Figure 1: The video frames generated by OmniHuman based on input audio and image. The generated results feature head and gesture movements, as well as facial expressions, that match the audio. OmniHuman generates highly realistic videos with any aspect ratio and body proportion, and significantly improves gesture generation and object interaction over existing methods, due to the data scaling up enabled by omni-conditions training.
footnotetext:  Equal contributionsfootnotetext:  Project lead
@@ 1 Introduction
Building upon these pretrained video diffusion networks, end-to-end human animation models, either for pose-driven human animation or audio-driven talking human generation, have developed rapidly since last year [18, 54, 70, 62, 8, 71, 52, 26, 34]. Despite achieving realistic results, these models are trained on highly filtered datasets to simplify the learning process, restricting their applicability to limited scenarios. For instance, most existing end-to-end audio-conditioned models are limited to facial or portrait animation, while most pose-conditioned models can only handle full-body images captured from a front-facing perspective with a static background. To date, no prior work has attempted to scale up training data for more generalizable human animation.
@@
Scaling up human animation data may seem straightforward, but unfortunately it is not. Directly adding more data is not always beneficial for network training. Take audio-conditioned models as an example: audio is primarily associated with facial expressions and has little correlation with body poses, background motion, camera movement, or lighting changes. As a result, raw training data must be filtered and cropped to minimize the influence of these unrelated factors. Additionally, audio-conditioned models often undergo further data cleaning based on lip-sync accuracy, which is also important to stabilize training. Similarly, pose-conditioned models require extensive filtering, cropping, and cleaning. Unfortunately, these processes discard a substantial amount of data, making dataset scaling a futile effort, despite the fact that much of the discarded data contains valuable motion patterns essential for training data expansion.
@@
In this paper, we address the challenges of scaling up human animation data and models. Our key insight is that incorporating multiple conditioning signals, such as text, audio, and pose, during training can significantly reduce data wastage. This approach offers two main advantages. On one hand, data that would otherwise be discarded for single-condition models (e.g., audio- or pose-conditioned) can be leveraged in tasks with weaker or more general conditions, such as text conditioning. Training on such data allows the model to learn more diverse motion patterns, mitigating the limitations imposed by data filtering. On the other hand, different conditioning signals can complement each other. For example, while audio alone cannot precisely control body poses, stronger conditions such as pose inputs can provide additional guidance. By integrating stronger conditioning signals alongside audio data during training, we aim to reduce overfitting and improve the generalization of generated results.
@@
Based on the above considerations, we designed the omni-conditions training strategy, which follows two proposed training principles: (1) stronger conditioned tasks can leverage weaker conditioned tasks and their corresponding data to achieve data scaling up during the model training process, and (2) the stronger the condition, the lower the training ratio that should be used. To implement this strategy, we built a mixed conditioned human video generation model named OmniHuman, based on the advanced video generation model architecture, DiT [42, 14]. OmniHuman can train with three motion-related conditions (text, audio, and pose) from weak to strong. This approach addresses the data scaling up challenge in end-to-end frameworks, allowing the model to benefit from large-scale data training, learn natural motion patterns, and support various input forms.
Overall, our contributions can be summarized as follows:
  1. 1.
    We propose the OmniHuman model, a mixed-conditioned human video generation model. It leverages our omni-conditions training strategy to integrate various motion-related conditions and their corresponding data. Unlike existing methods that reduce data due to stringent filtering, our approach benefits from large-scale mixed conditioned data.
  2. 2.
    OmniHuman generates highly realistic and vivid human motion videos, supporting multiple modalities simultaneously. It performs well with different portrait and input aspect ratios. OmniHuman significantly improves gesture generation, a challenge for previous end-to-end models, and supports various image styles, significantly outperforming existing audio-conditioned human video generation methods.
@@ 2.2 Human Animation
Figure 2: The framework of OmniHuman. It consists of two parts: (1) the OmniHuman model, which is based on the DiT architecture and supports simultaneous conditioning with multiple modalities including text, image, audio, and pose; (2) the omni-conditions training strategy, which employs progressive, multi-stage training based on the motion-related extent of the conditions. The mixed condition training allows the OmniHuman model to benefit from the scaling up of mixed data.
@@ 3 Method
In this section, we introduce our framework, OmniHuman, which employs motion-related condition mixing during network training to scale up the training data. First, we provide an overview of the framework, including its inputs, outputs and key design elements. Next, we focus on the omni-conditions design, covering audio, pose, and reference conditions. We then detail the training strategy of OmniHuman, which leverages these omni-conditions for mixed data training, enabling the model to learn natural motion from large-scale datasets. Finally, we describe the implementation details for the inference phases of the OmniHuman model.
@@ 3.1 Overview
As illustrated in Figure 2, our approach consists of two primary parts: the OmniHuman model, a multi-condition diffusion model and the Omni-Conditions Training Strategy. For model, The OmniHuman model begins with a pretrained Seaweed model [35], which uses MMDiT [14, 42] and is initially trained on general text-video pairs for text-to-video and text-to-image tasks. Given a reference image, the OmniHuman model aims to generate human videos using one or more driving signals including text, audio and pose. To achieve this, we employ various strategies to integrate frame-level audio features and pose heatmap features into the OmniHuman model. The detailed procedure is explained in the following subsections. OmniHuman model utilizes a causal 3DVAE [80] to project videos at their native size [12] into a latent space and employs flow matching [36] as the training objective to learn the video denoising process. We employ a three-stage mixed condition post-training approach to progressively transform the diffusion model from a general text-to-video model to a multi-condition human video generation model. As depicted on the left of Figure 2, these stages sequentially introduce the driving modalities of text, audio, and pose according to their motion correlation strength, from weak to strong, and balance their training ratios.
3.2 Omni-Conditions Designs
@@
Driving Conditions. We adopted different approaches for injecting audio and pose conditions. Regarding audio condition, the wav2vec [45, 1] model is employed to extract acoustic features, which are subsequently compressed using a MLP to align with the hidden size of MMDiT. The features of each frame are concatenated with the audio features from adjacent timestamps to generate audio tokens for the current frame. As depicted in Figure 2, these audio tokens are injected into each block of MMDiT through cross-attention, enabling interaction between the audio tokens and the noisy latent representations. To incorporate pose condition, we use a pose guider to encode the driving pose heatmap sequence. The resulting pose features are concatenated with those of adjacent frames to acquire pose tokens. These pose tokens are then stacked with the noise latent along the channel dimension and fed into the unified multi-condition diffusion model for visual alignment and dynamic modeling. The text condition is retained as in the MMDiT text branch.
@@
Appearance Conditions. The goal of OmniHuman is to generate video outputs that preserve both the subject’s identity and the background details from a reference image. To achieve this, previous research has proposed various strategies for injecting appearance representations into the denoising process. The most widely adopted approach involves using a reference network [26, 34, 54], a parallel, trainable copy of the entire diffusion UNet or DiT that integrates with the self-attention layers of the original denoising Net. While effective at transferring appearance features to the denoising process, this method requires duplicating a full set of trainable parameters, which presents scalability challenges as model size increases. To overcome this challenge, OmniHuman introduces a simple yet effective strategy for reference conditioning. Instead of constructing additional network modules, we reuse the original denoising DiT backbone to encode the reference image. Specifically, the reference image is first encoded into a latent representation using a VAE, and both the reference and noisy video latents are flattened into token sequences. These sequences are then packed together and simultaneously fed into the DiT, enabling the reference and video tokens to interact via self-attention across the entire network. To help the network distinguish between reference and video tokens, we modify the 3D Rotational Position Embeddings (RoPE) [53] in the DiT by zeroing the temporal component for reference tokens, while leaving the RoPE for video tokens unchanged. This approach effectively incorporates appearance conditioning without adding extra parameters. In addition to the reference image, to support long video generation, we draw on previous methods by using motion frames [52], concatenating their features with the noise features.
After introducing these conditions, the motion-related conditions now include text, reference image, audio, and pose. Text describes the current event, the reference image defines the range of motion, audio determines the rhythm of co-speech gestures, and pose specifies the exact motion. Their correlation strength with human motions can be considered to decrease in this order.
3.3 Scaling up with Omni-Conditions Training
Thanks to the multi-condition design, we can divide the model training into multiple tasks, including image and text to video, image and text, audio to video, and image and text, audio, pose to video. During training, different modalities are activated for different data, allowing a broader range of data to participate in the training process and enhancing the model’s generation capabilities. After the conventional text-to-video pretraining phase, we follow two training principles for scaling up the conditioned human video generation task. Principle 1, stronger conditioned tasks can leverage weaker conditioned tasks and their corresponding data to achieve data scaling up during the model training process. Data excluded from audio and pose conditioned tasks due to filtering criteria like lip-sync accuracy, pose visibility, and stability can be used in text and image conditioned tasks, as they meet the standards for weaker conditions. Therefore, in the first stage 1, we drop the audio and pose conditions. Principle 2, the stronger the condition, the lower the training ratio that should be used. During training, stronger motion-related conditions, such as pose, generally train better than weaker conditions like audio due to less ambiguity. When both conditions are present, the model tends to rely on the stronger condition for motion generation, preventing the weaker condition from learning effectively. Therefore, we ensure that weaker conditions have a higher training ratio than stronger conditions. We construct stage 2 to drop only the pose condition, and in the final stage 3, use all conditions. Additionally, the training ratios for text, reference, audio, and pose are progressively halved. This approach assigns higher gradient weights to more challenging tasks and prevents overfitting to a single condition during overlapping condition training. Principle 1 allows us to significantly expand the training data, while Principle 2 ensures that the model fully utilizes the advantages of each motion-related condition during mixed conditions training and learns their motion generation capabilities. By combining Principles 1 and 2, OmniHuman can effectively train with mixed conditioned data, benefiting from data scaling up and achieving satisfactory results.
@@ 3.4 Inference Strategies
For audio-driven scenarios, all conditions except pose are activated. For pose-related combinations, all conditions are activated, but for pose-only driving, audio is disabled. Generally, when a condition is activated, all conditions with a lower motion-related influence are also activated unless unnecessary. During inference, to balance expressiveness and computational efficiency, we apply classifier-free guidance (CFG) [20] specifically to audio and text across multiple conditions. However, we observed that an increased CFG results in pronounced wrinkles on the characters, whereas a decreased CFG compromises lip synchronization and motion expressiveness. To mitigate these issues, we propose a CFG annealing strategy that progressively reduces the CFG magnitude throughout the inference process, thereby significantly minimizing the appearance of wrinkles while ensuring that expressiveness. OmniHuman is capable of producing video segments of arbitrary length within memory constraints based on the provided reference images and various driving signals. To ensure temporal coherence and identity consistency in long videos, the last five frames of the previous segment are utilized as motion frames.

### v3

@@ Abstract
End-to-end human animation, such as audio-driven talking human generation, has undergone notable advancements in the recent few years. However, existing methods still struggle to scale up as large general video generation models, limiting their potential in real applications. In this paper, we propose OmniHuman, a Diffusion Transformer-based framework that scales up data by mixing motion-related conditions into the training phase. To this end, we introduce two training principles for these mixed conditions, along with the corresponding model architecture and inference strategy. These designs enable OmniHuman to fully leverage data-driven motion generation, ultimately achieving highly realistic human video generation. More importantly, OmniHuman supports various portrait contents (face close-up, portrait, half-body, full-body), supports both talking and singing, handles human-object interactions and challenging body poses, and accommodates different image styles. Compared to existing end-to-end audio-driven methods, OmniHuman not only produces more realistic videos, but also offers greater flexibility in inputs. It also supports multiple driving modalities (audio-driven, video-driven and combined driving signals).
Figure 1: The video frames generated by OmniHuman based on input audio and image. The generated results feature head and gesture movements, as well as facial expressions, that match the audio. OmniHuman generates realistic videos with any aspect ratio and body proportion, and significantly improves gesture generation and object interaction over existing methods, due to the data scaling up enabled by omni-conditions training.
footnotetext: Equal contributionsfootnotetext: Project lead and corresponding author: jianwen.alan@gmail.com
@@ 1 Introduction
Building upon these pretrained video diffusion networks, end-to-end human animation models, especially for audio-driven talking human generation, have developed rapidly since last year [19, 54, 68, 60, 9, 69, 51, 26, 34]. Despite achieving realistic results, these models are trained on highly filtered datasets to simplify the learning process, restricting their applicability to limited scenarios. For instance, most existing end-to-end audio-driven models are limited to facial or portrait images captured from a front-facing perspective with a static background. To date, no prior work has attempted to scale up training data for more generalizable human animation.
@@
Scaling up human animation data may seem straightforward, but unfortunately it is not. Directly adding more data is not always beneficial for network training. Take audio-conditioned models as an example: audio is primarily associated with facial expressions and has little correlation with body poses, background motion, camera movement, or lighting changes. As a result, raw training data must be filtered and cropped to minimize the influence of these unrelated factors. Additionally, audio-conditioned models often undergo further data cleaning based on lip-sync accuracy, which is also important to stabilize training. Unfortunately, these processes discard a substantial amount of data, making dataset scaling a futile effort or difficult to achieve, despite much of the discarded data containing valuable motion patterns essential for training data expansion. For example, recent state-of-the-art methods mention that after rigorous cleaning, less than 10% of the data is retained, rendering direct data scaling highly cost-ineffective.
@@
In this paper, we address the challenges of scaling up audio-driven human animation data and models. Our key insight is that incorporating multiple conditioning signals beyond audios, such as text and pose, during training can significantly reduce data wastage. This approach offers two main advantages. On one hand, data that would otherwise be discarded for single-condition models (e.g., audio-driven) can be leveraged in tasks with weaker or more general conditions, such as text. Training on such data allows the model to learn more diverse motion patterns, mitigating the limitations imposed by data filtering. On the other hand, different conditioning signals can complement each other. For example, while audio alone cannot precisely control body poses, stronger conditions such as body poses can provide additional guidance. By integrating stronger conditioning signals alongside audio data, we aim to reduce overfitting and improve the generalization of generated results.
@@
Building on the above considerations, we designed an omni-conditions training strategy, which follows two key training principles: (1) tasks with stronger conditioning can leverage those with weaker conditioning, along with their associated data, to scale up training, and (2) the stronger the conditioning signal, the lower its training ratio should be. To implement these strategies, we develop OmniHuman, a mixed conditioned human video generation model based on the advanced video generation architecture, DiT [41, 15]. Although OmniHuman mainly focuses on audio-driven human video generation, it is trained with three motion-related conditions, including text, audio, and pose, ranging from weak to strong. This approach effectively addresses the challenge of data scaling in end-to-end audio-driven frameworks, allowing the model to learn natural motion patterns from large-scale data and support various input forms.
In summary, we introduce OmniHuman, an audio-driven human video generation model that leverages our omni-conditions training strategy to integrate various motion-related conditions and their corresponding data. Unlike existing methods that reduce data due to strict filtering, our approach benefits from large-scale mixed-conditioned data, enabling OmniHuman to generate highly realistic and expressive human motion videos given a reference image and the driving audio. It adapts to various portrait types and aspect ratios, and significantly improves gesture generation, which is a longstanding challenge for previous methods. Additionally, it accommodates diverse image styles and background contents. To the best of our knowledge, OmniHuman is the first solution capable of audio-driven human video generation on input images with any body proportions and image styles, and also supports auxiliary pose driving.
@@ 2.2 Human Animation
Figure 2: The framework of OmniHuman. It consists of two parts: (1) the OmniHuman model, which is based on the DiT architecture and supports simultaneous conditioning with multiple modalities including text, image, audio, and pose. To support long video continuation, we concatenate the latents of the last generated frames with noise latents, which are omitted for simplicity. (2) the omni-conditions training strategy, which employs progressive, multi-stage training based on the motion-related extent of the conditions. The mixed condition training allows the OmniHuman model to benefit from the scaling up of mixed data.
@@ 3.1 Overview
The goal of OmniHuman is to generate realistic human animation videos from a reference image and using one or more driving conditions, including text, audio and pose. Here we mainly focuses on audio-conditioned generation, and leverage the other two as auxiliary signals. To achieve this, our approach consists of two primary parts: the OmniHuman model, a multi-condition diffusion model, and the Omni-conditions Training Strategy. Figure 2 illustrates our approach. Our OmniHuman model is built upon an MMDiT [15, 41, 35, 13]-based text-to-video model, which is initially trained on general text-video pairs. We design different methods to integrate different driving signals, including audio features and skeleton map features, into the OmniHuman model (Sec. 3.2). Then we employ a three-stage mixed condition training approach to progressively transform the diffusion model from a general text-to-video model to a multi-condition human video generation model (Sec. 3.3). As illustrated on the left of Figure 2, these stages sequentially introduce the driving modalities of text, audio, and pose based on their motion correlation strength, from weak to strong, and balance their training ratios. This ultimately enables the model to be compatible with different modalities and benefit from large-scale data with mixed conditions.
Note that OmniHuman operates in a compact latent space by using a causal 3D VAE [78] to project videos at their native resolution into compact latents. However, for simplicity, we omit this process in the following sections.
3.2 Omni-Conditions Module Design
For designing the condition injection module for OmniHuman, we adhere to a minimalist design principle, aiming to minimize modality-specific parameters. This ensures that multimodal interactions can be more extensively modeled within the shared MMDiT backbone, which contains the majority of model parameters. We categorize the roles of modalities in the human animation model into two types of conditions, as discussed below.
@@ 3.2 Omni-Conditions Module Design
Driving Conditions. In OmniHuman, driving conditions include audio, pose, and text, with audio being the primary focus of this paper. For the audio condition, the wav2vec model [44, 1] is employed to extract multi-scale acoustic features [26, 54]. These features are subsequently compressed using an MLP to align with the hidden size of MMDiT and the framerate of videos (25 fps in our setting). The features of each frame are concatenated with audio features from adjacent timestamps to generate audio tokens for the current frame. As depicted in Figure 2, these audio tokens are injected into each block of MMDiT through frame-wise cross-attention, enabling interaction between the audio tokens and the noisy latent representations.
To incorporate the pose condition, we use a pose guider [25] to encode the driving skeleton map sequence. The resulting pixel-aligned pose features are concatenated with those of adjacent frames to acquire frame-wise pose tokens. These pose tokens are then stacked with the noisy latent representations along the channel dimension and fed into the model for visual alignment and dynamic modeling.
The text condition is processed with the original MMDiT’s text branch, following the common practice of text-to-image and text-to-video frameworks [15, 41].
@@
Appearance Conditions. The appearance conditioning is incorporated to preserve the identity of the foreground character while ensuring natural background motion. Previous research often used a reference network approach [26, 34, 54], which involves a parallel, trainable copy of the entire diffusion backbone that integrates with the self-attention layers of the original denoising network. While effective, this method requires duplicating a full set of trainable parameters, posing scalability challenges as model size increases. In contrast, our OmniHuman adopts a simpler yet effective strategy for reference conditioning by reusing the original denoising DiT backbone to encode the reference image. Specifically, both the reference and noisy video latents are flattened into token sequences, packed together, and simultaneously fed into the DiT. This allows the reference and video tokens to interact via self-attention across the entire network. To distinguish between reference and video tokens, we modify the 3D Rotational Position Embeddings (RoPE) [52] in the DiT by zeroing the temporal component for reference tokens while leaving the RoPE for video tokens unchanged. Such an approach effectively incorporates appearance conditioning without adding extra parameters, maximizing the modeling of interactions between different modalities within the shared parameters.
In addition to the reference image, to support long video generation, we use motion frames [51] by simply concatenating their features with the noise tokens as in [78].
3.3 Scaling up with Omni-Condition Training
Thanks to its multi-condition design, OmniHuman enables training across multiple tasks, including: (1) image and text to video, (2) image, text, and audio to video, and (3) image, text, audio, and pose to video. During training, different modalities are activated based on the available data, allowing a wider range of data to contribute to the training process and enhancing the model’s generative capabilities. Starting with a pretrained general text-to-video model, we follow two key training principles to scale up conditioned human video generation, progressively transforming the pretrained model into a mixed-conditioned human generation framework.
Principle 1: Stronger conditioned tasks can leverage weaker conditioned tasks and their corresponding data to scale up training data during model training. For instance, data excluded from audio-conditioned tasks due to filtering criteria, such as lip-sync accuracy and stability, can still be utilized in text- and image-conditioned tasks, as they meet the standards for weaker conditions. Therefore, in Stage 1, we exclude the audio and pose conditions and train the model on the text+image-to-video task, maximizing the use of available training data.
Principle 2: The stronger the condition, the lower its training ratio should be. During training, stronger motion-related conditions, such as pose, generally provide more precise motion control compared to weaker conditions like audio, which may introduce ambiguity. When both conditions are present, the model tends to rely on the stronger condition, limiting the learning effectiveness of the weaker one. To achieve a unified multi-condition human animation model, we balance this effect by assigning a higher training ratio to weaker conditions and adjusting their training order, thus ensuring they contribute meaningfully to motion generation. To be more specific, we drop only the pose condition in Stage 2 and incorporate all conditions in the final Stage 3. The training ratios for text, audio, and pose are progressively halved when present, assigning higher gradient weights to more challenging tasks and preventing overfitting to a single condition during mixed-condition training.
In summary, Principle 1 allows us to significantly expand the training data, while Principle 2 ensures that the model fully leverages each motion-related condition’s strengths during mixed-condition training, enabling it to learn distinct motion generation capabilities and avoiding either overfitting or underfitting. By integrating both principles, OmniHuman effectively trains on mixed-conditioned data, benefiting from data scaling while achieving high-quality results.
@@ 3.4 Inference Strategies
For audio-driven scenarios, all conditions except pose are activated. For mixed driving with pose and audio, all conditions are activated. For pose-only driving, audio is disabled. In general, when a condition is activated, all conditions with a lower motion-related influence are also activated unless deemed unnecessary. Additionally, for all animation tasks, we use an image captioning model to obtain the text input.
During inference, to balance expressiveness and computational efficiency, we apply classifier-free guidance (CFG) specifically to audio and text, but not to the pose condition. The CFG scale is set to 6.5. We found that only conditions with relatively weaker motion-related influence, such as audio and text, require CFG. To maintain temporal coherence and identity consistency in long videos, the last five frames of the previous segment are used as motion frames for the following segment.
Table 1: Quantitative analysis of Omni-Conditions Training. The upper and lower parts correspond to Principles 1 and 2 respectively.
Methods CelebV-HQ RAVDESS CyberHost
[…]
Ref Img 3.195 1.963 - - - 4.421 2.578 - - - 3.963 2.634 - - - - -
$0\%$ T-Data 3.855 2.693 4.299 39.80 47.86 4.671 3.053 4.663 18.20 18.83 4.262 3.171 6.465 38.38 45.77 35.82 0.871
$25\%$ T-Data 3.758 2.571 3.311 37.95 47.04 4.652 3.011 3.552 17.03 16.00 4.236 3.152 5.039 38.36 43.69 40.39 0.877
$50\%$ T-Data 3.717 2.525 3.696 36.26 46.22 4.541 2.913 3.826 16.92 15.51 4.217 3.132 5.591 37.06 37.95 40.69 0.872
$100\%$ T-Data 3.644 2.452 4.987 36.01 43.74 4.445 2.763 5.092 16.70 15.13 4.076 2.959 6.938 33.27 31.15 43.54 0.882
IA 3.644 2.452 4.987 36.01 43.74 4.445 2.763 5.092 16.70 15.13 4.076 2.959 6.938 33.27 31.15 43.54 0.882
IPA 3.458 2.279 2.788 38.98 44.70 4.275 2.681 2.923 21.20 25.05 3.945 2.815 4.879 39.60 36.03 45.44 0.822
IAP, A<P 3.556 2.370 4.201 38.73 44.63 4.298 2.689 4.752 20.79 17.18 3.979 2.869 6.765 34.26 32.79 40.99 0.869
IAP, A>P 3.602 2.415 4.934 36.66 43.36 4.425 2.796 4.882 17.43 15.66 4.055 2.959 6.951 33.27 31.08 39.39 0.886
4 Experiments
4.1 Implementation Details
Dataset. We use 18.7K hours of in-house human-related data for training, which are filtered based on aesthetics, image quality, motion amplitude, etc. (common criteria for video generation [8]). Of this dataset, 13% was selected using lipsync and pose visibility criteria, enabling audio and pose modalities. During training, the data composition was adjusted to fit the omni-condition training strategy. For testing, we conduct the evaluation following the portrait animation method Loopy [26] and the half-body animation method CyberHost [34]. We randomly sampled 100 videos from public portrait datasets, including CelebV-HQ [81] (a diverse dataset with mixed scenes) and RAVDESS [28] (an indoor dataset including speech and song) as the testset for portrait animation. For half-body animation, we used CyberHost’s test set, which includes a total of 269 body videos with 119 identities, encompassing different races, ages, genders, and initial poses.
During training, we use a learning rate of $5\times 10^{-5}$ with the AdamW optimizer, a gradient clip of 1.0, and a batch size of 256. The weight decay is set to 0.01. The training process is carried out on 400 A100 GPUs, with each phase lasting approximately 10 days.
In the following sections, we first validate the effectiveness of the OmniHuman design through ablation studies, primarily using audio driving for comparison because it not only reflects mimicking ability but also better showcases motion generation capability. For efficiency, experiments are conducted at 480P resolution. Next, we present several visual results and applications to illustrate the generalization capability of OmniHuman. Finally, we compare our results at 720P resolution with current state-of-the-art methods to demonstrate OmniHuman’s versatility across different tasks. It is important to note that our aim is to demonstrate that the OmniHuman paradigm better facilitates data scaling for more general human animation rather than focusing on performance in limited scenarios. This could potentially be an important training paradigm for human animation methods.